Shifter leader was rk Shifters were nt 1. The beam was running from midnight through 5:15am, and we lost the beam unexpectedly. New beam started around 6am and continuing to next shift (8am through 4pm on Monday June, 12, 2006). 2. We run 12 times from run #16078 - #16092. FS at 3 degrees 1723 A and MRS at 45 degrees 1050 B. 3. T3 tripped once at 3:15am. _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Mon Jun 12 08:00:02 2006
This archive was generated by hypermail 2.1.8 : Mon Jun 12 2006 - 08:00:10 EDT